# Zihan Zhang Princeton

### v4

@@ 6 Necessity of randomization
  • a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$,
@@
In words, the sample complexity when a deterministic output is required scales at least with $\Omega(\frac{dk}{\varepsilon^{2}}\big{)}$, which is considerably larger than the sample complexity $\widetilde{O}\big{(}\frac{d+k}{\varepsilon^{2}}\big{)}$ achievable via Algorithm 1 (which outputs a randomized hypothesis). The proof of this theorem is deferred to Appendix D.
@@ 7.1 Preliminaries: Rademacher complexity
Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in, e.g., shalev2014understanding.
@@
For instance, if the hypothesis class $\mathcal{H}$ obeys $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, then it is well-known that Assumption 1 holds with the choice (see, e.g., mohri2018foundations)
@@
One might raise a natural question about Assumption 1: can we use $\widetilde{C}_{n}\coloneqq\max_{i\in[k]}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ instead of ${C}_{n}$ without incuring a worse sample complexity? The answer is, however, negative. In fact, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, and hence we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$. The interested reader is referred to Appendix E.4 for more details.
@@ 7.2 Algorithm and sample complexity
       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
@@ 10                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
In view of Lemma 2 and Lemma 3, it boils down to showing that running Algorithm 3 results in $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$, a property that holds with probability at least $41-\delta/$ 4. To accomplish this, we have the lemma below.
@@
Suppose Assumption 1 holds. With probability at least $41-\delta/$ 4, the iterates of Algorithm 3 satisfy
@@
The proof of Lemma 6 is postponed to Appendix E. Combine this with Lemma 2 and Lemma 3 to show that: the total number of samples collected in Algorithm 3 is upper bounded by
@@ 8 Discussion
In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming availability of adaptive (or on-demand) sampling. We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor. From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex-concave optimization problem. We have further unveiled the necessesity of randomization, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses. Notably, our work manages to solve three open problems presented in COLT 2023 (namely, awasthi2023sample).
Our work not only addresses existing challenges but also opens up several directions for future exploration. To begin with, while our sample complexity results are optimal up to logarithmic factors, further studies are needed in order to sharpen the logarithmic depdency. Additionally, the current paper assumes a flexible sampling protocol that allows the learner to take samples arbitrarily from any of the $k$ distributions; how will the sample complexity be impacted under additional constraints imposed on the sampling process? Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning? Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (hebert2018multicalibration; haghtalab2023unifying). We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
@@ Acknowledgements
We thank Eric Zhao for answering numerous questions about the open problems. YC is supported in part by the Alfred P. Sloan Research Fellowship, the NSF grants CCF-1907661, DMS-2014279, IIS-2218713 and IIS-2218773. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, the Sloan Research Fellowship, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994.
@@ Appendix A Auxiliary lemmas
We begin by introducing three handy concentrations inequalities. The first result is the well-renowned Freedman inequality (freedman1975tail), which assists in deriving variance-aware concentration inequalities for martingales.
Lemma 7 (Freedman’s inequality (freedman1975tail)).
@@
Lemma 8 (Lemma 10 in zhang2022horizon).
@@
Clearly, it suffices to prove (44) for the special case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fixing $\sigma_{i}$ for $11\leq i\leq n-$ 1, we can deduce that
@@ B.1 Proof of Lemma 1
Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
@@
    =$\mathcal{B}=\bigg{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\,\Big{|}\,\frac{n_{i}}{w_{i}}\geq\frac{T_{1}}{2},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\bigg{\}},$
@@
one can see from (54) that, with probability exceeding $41-\delta/$ 4,
@@ B.2 Proof of Lemma 2
Before proceeding, let us introduce some additional notation. Let $\delta^{\prime}\coloneqq\frac{\delta}{4(T+k+1)}$, and define
$\mathsf{OPT}\coloneqq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})$
@@
$v^{t}\leq\min_{h\in\mathcal{H}}L(h,w^{t})-\mathsf{OPT}+\varepsilon_{1}=\min_{h\in\mathcal{H}}L(h,w^{t})-\min_{h\in\mathcal{H}}\max_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon_{1}\leq\varepsilon_{1},\qquad\forall 1\leq t\leq T.$ (59)
@@
[…]

### v5

@@ 6 Limitations of proper learning?
  • a loss function $\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$,
@@
In words, the sample complexity when a deterministic output from $\mathcal{H}$ is required scales at least with $\Omega(\frac{dk}{\varepsilon^{2}}\big{)}$, which is considerably larger than the sample complexity $\widetilde{O}\big{(}\frac{d+k}{\varepsilon^{2}}\big{)}$ achievable via Algorithm 1 (which outputs a randomized hypothesis). This demonstrates the benefits of randomization, or more broadly, the necessity of improper learning. The proof of this theorem is deferred to Appendix C.
@@ 7.1 Preliminaries: Rademacher complexity
Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in, e.g., Shalev-Shwartz and Ben-David, (2014).
@@
For instance, if the hypothesis class $\mathcal{H}$ obeys $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, then it is well-known that Assumption 1 holds with the choice (see, e.g., Mohri et al., (2018))
@@
One might raise a natural question about Assumption 1: can we use $\widetilde{C}_{n}\coloneqq\max_{i\in[k]}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ instead of ${C}_{n}$ without incuring a worse sample complexity? The answer is, however, negative. In fact, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, and hence we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$. The interested reader is referred to Appendix D.4 for more details.
@@ 7.2 Algorithm and sample complexity
  /* recompute $\widehat{w}^{t}$ & draw new samples for $\mathcal{S}_{\mathrm{w}}$ only if $w^{t}$ changes sufficiently. */
@@ 10        $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
In view of Lemma 2 and Lemma 3, it boils down to showing that running Algorithm 3 results in $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$, a property that holds with probability at least $1-\delta/4$. To accomplish this, we have the lemma below.
@@
Suppose Assumption 1 holds. With probability at least $1-\delta/4$, the iterates of Algorithm 3 satisfy
@@
The proof of Lemma 11 is postponed to Appendix D. Combine this with Lemma 2 and Lemma 3 to show that: the total number of samples collected in Algorithm 3 is upper bounded by
@@
8 Extension: multi-loss multi-distribution learning
In this section, we consider an extension of the MDL problem to the multi-loss (or multi-objective) setting.
Problem setting. Formally, consider again the hypothesis class $\mathcal{H}$ and the $k$ distributions of interest $\{\mathcal{D}_{i}\}_{i=1}^{k}$. Suppose that there are $R\geq 2$ loss functions $\mathcal{L}\coloneqq\{\ell^{j}\}_{j=1}^{R}$ to consider, where $\ell^{j}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\rightarrow[-1,1]$ denotes the $j$-th loss function of interest. The goal is to find a (possibly randomized) hypothesis $\widehat{h}\in\mathcal{H}$ such that
$\max_{1\leq i\leq k}\max_{\ell\in\mathcal{L}}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},\widehat{h}}\big{[}\ell\big{(}\widehat{h},(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k,\ell\in\mathcal{L}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (58)
Evidently, this formulation is a natural extension of the single-loss MDL problem. It is particularly relevant in scenarios where multiple criteria need to be satisfied or optimized simultaneously, such as those involving fairness and robustness. It has also been shown that this extension could be helpful towards solving the multicalibration problem (Haghtalab et al.,, 2023).
Additional notation. Let us introduce one more notation to be used throughout. For any $i\in[k]$ and $\ell\in\mathcal{L}$, we denote by $L^{\ell}_{i}(h)$ the expected loss of $h$ w.r.t. $\mathcal{D}_{i}$ and $\ell$, that is,
$L_{i}^{\ell}(h)\coloneqq\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}.$ (59)
Then for any $u\in\Delta(\mathcal{D}\times\mathcal{L})$, define
$L(h,u)\coloneqq\sum_{i,\ell}u_{i,\ell}L^{\ell}_{i}(h).$ (60)
The proposed algorithm and main theorem. As it turns out, Algorithm 1 can be extended to tackle the above multi-loss multi-distribution learning problem. The full algorithm is presented in Algorithm 4. As a key distinction from Algorithm 1, we need to maintain a Hedge-type algorithm over $\mathcal{D}\times\mathcal{L}$ in an attempt to solve the following problem:
$\max_{u\in\Delta(\mathcal{D}\times\mathcal{L})}\min_{h\in\mathcal{H}}L(h,u).$
Akin to Algorithm 1, in each round $t$, we maintain a dataset $\mathcal{S}$ to construct estimation of $L(h,u^{t})$ for all $h\in\mathcal{H}$, which help us to compute the optimal response $h^{t}$ with respect to the current weight $u^{t}$. Then we query a small number (at most $\tilde{O}(k)$) of new samples to assist in estimating each $L^{\ell}_{i}(h^{t})$. Notably, we need to choose $T_{1}$ to be larger than the case of Algorithm 1, in order to apply the union bound argument. Given that $u^{t}$ is in the $kR$-dimensional simplex $\Delta(\mathcal{D}\times\mathcal{L})$, applying the union bound uniformly over all possible choices of $u^{t}$ would lead to an undesirable factor of $\widetilde{O}(kR)$. Instead of this analysis strategy, we consider controlling an upper bound on the error , which allows us to pay a smaller factor of $O(k\log(R))$ instead.
The following theorem establishes the sample complexity of Algorithm 4.
Theorem 4.
There exists an algorithm (see Algorithm 4 for details) such that: with probability exceeding $1-\delta$, the randomized hypothesis $h^{\mathsf{final}}$ returned by this algorithm achieves
$\max_{1\leq i\leq k}\max_{\ell\in\mathcal{L}}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h^{\mathsf{final}}}\big{[}\ell\big{(}h^{\mathsf{final}},(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k,\ell\in\mathcal{L}}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon,$
provided that the total sample size exceeds
$\frac{\big{(}d+k\log(R)\big{)}\min\{\log(R),k\}}{\varepsilon^{2}}\mathrm{poly}\log\left(k,d,\frac{1}{\varepsilon},\frac{1}{\delta},\log(R)\right).$ (61)
The full proof of Theorem 4 is provided in Appendix E. Most parts of the proof are identical to that of Theorem 1, except that: (a) the new method in estimating $\{L(h,u^{t})\}_{h\in\mathcal{H}}$ mentioned above (see also Lemma 22); (b) a more refined analysis of the trajectory bound to remove some $\mathrm{poly}(\log(R))$ factors (see Lemma 24).
In comparison to the single-loss case (see Theorem 1), the multi-loss version of our algorithm only incurs an additional price of $\mathrm{poly}(\log(R))$ factors, which is particularly appealing when the number $R$ of loss functions is no greater than a polynomial function in $d,k,$ etc.
1input: $k$ data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, loss function class $\mathcal{L}=\{\ell^{j}\}_{j=1}^{R}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$.
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{kR}{\delta\varepsilon}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}:=\frac{40000\left(k\log(\frac{kR}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}.$
3 initialization: for all $i\in[k]$ and $\ell\in\mathcal{L}$, set $U^{1}_{i,\ell}=1$, $\widehat{w}^{0}_{i}=0$ and $n_{i}^{0}=0$; $\mathcal{S}=\emptyset$.
4 for $t=1,2,\ldots,T$ do
5   set $u^{t}=[u^{t}_{i,\ell}]_{1\leq i\leq k,\ell\in\mathcal{L}}$ with $u^{t}_{i,\ell}\leftarrow\frac{U^{t}_{i,\ell}}{\sum_{i^{\prime}\in[k],\ell^{\prime}\in\mathcal{L}}U^{t}_{i^{\prime},\ell^{\prime}}}$ .
6   set $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$ with $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
  /* recompute $\widehat{w}^{t}$ & draw new samples for $\mathcal{S}$ only if $w^{t}$ changes sufficiently. */
7 if there exists $j\in[k]$ such that $\sum_{\ell\in\mathcal{L}}u_{j,\ell}^{t}\geq 2\hat{w}_{j}^{t-1}$  then
8    $\widehat{w}^{t}_{i}\leftarrow\max\big{\{}\sum_{\ell\in\mathcal{L}}u_{i,\ell}^{t},\widehat{w}^{t-1}_{i}\big{\}}$ for all $i\in[k]$;
9    for $i=1,\ldots,k$ do
10       $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
12
13   compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,u^{t})$, where
$\widehat{L}^{t}(h,u^{t}):=\sum_{i=1}^{k}\sum_{\ell\in\mathcal{L}}\frac{u^{t}_{i,\ell}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (62)
  /* estimate the loss vector and execute weighted updates. */
14 $\overline{w}^{t}_{i}\leftarrow\max_{1\leq\tau\leq t}\sum_{\ell\in\mathcal{L}}u^{\tau}_{i,\ell}$ for all $i\in[k]$.
15 for $i=1,\ldots,k$ do
[…]
    update the weight as $U^{t+1}_{i,\ell}=U^{t}_{i,\ell}\exp(\eta\widehat{r}_{i,\ell}^{t})$. // Hedge updates.
17
18
Algorithm 4 Hedge for multi-loss multi-distribution learning ($\mathtt{MLMDL}\text{-}\mathtt{Hedge}\text{-}\mathtt{VC}$)
@@ 9 Discussion
In this paper, we have settled the problem of achieving optimal sample complexity in multi-distribution learning, assuming availability of adaptive (or on-demand) sampling. We have put forward a novel oracle-efficient algorithm that provably attains a sample complexity of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ for VC classes, which matches the best-known lower bound up to some logarithmic factor. From the technical perspective, the key novelty of our analysis lies in carefully bounding the trajectory of the Hedge algorithm on a convex-concave optimization problem. We have further unveiled the necessity of improper learning, revealing that a considerably larger sample size is necessary if the learning algorithm is constrained to return deterministic hypotheses from $\mathcal{H}$. Notably, our work manages to solve three open problems presented in COLT 2023 (namely, Awasthi et al., (2023, Problems 1, 3 and 4)).
Our work not only addresses existing challenges but also opens up several directions for future exploration. To begin with, while our sample complexity results are optimal up to logarithmic factors, further studies are needed in order to sharpen the logarithmic dependency. Additionally, the current paper assumes a flexible sampling protocol that allows the learner to take samples arbitrarily from any of the $k$ distributions; how will the sample complexity be impacted under additional constraints imposed on the sampling process? Furthermore, can we extend our current analysis (which bounds the dynamics of the Hedge algorithm) to control the trajectory of more general first-order/second-order algorithms, in the context of robust online learning? Another venue for exploration is the extension of our multi-distribution learning framework to tackle other related tasks like multi-calibration (Hébert-Johnson et al.,, 2018; Haghtalab et al.,, 2023). We believe that our algorithmic and analysis framework can shed light on making progress in all of these directions.
@@ Acknowledgements
We thank Eric Zhao for answering numerous questions about the open problems. We would also like to thank Chicheng Zhang for pointing out an error in a preliminary version of this paper. YC is supported in part by the Alfred P. Sloan Research Fellowship, the NSF grants CCF-1907661, DMS-2014279, IIS-2218713 and IIS-2218773. JDL acknowledges support of the ARO under MURI Award W911NF-11-1-0304, the Sloan Research Fellowship, NSF CCF 2002272, NSF IIS 2107304, NSF CIF 2212262, ONR Young Investigator Award, and NSF CAREER Award 2144994.
@@ Appendix A Auxiliary lemmas
We begin by introducing three handy concentrations inequalities. The first result is the well-renowned Freedman inequality (Freedman,, 1975), which assists in deriving variance-aware concentration inequalities for martingales.
Lemma 12 (Freedman’s inequality (Freedman,, 1975)).
@@
Lemma 13 (Lemma 10 in Zhang et al., (2022)).
@@
Clearly, it suffices to prove (65) for the special case where $w_{i}^{1}=w^{2}_{i}$ for $1\leq i\leq n-1$, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fixing $\sigma_{i}$ for $1\leq i\leq n-1$, we can deduce that
@@ B.1 Proof of Lemma 1
Note that for any dataset $\mathcal{C}$ with cardinality $|\mathcal{C}|$, the Sauer–Shelah lemma (Wainwright,, 2019, Proposition 4.18) together with our assumption that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ tells us that the cardinality of the following set obeys
@@
    $\mathcal{B}=\bigg{\{}n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\,\Big{|}\,\frac{n_{i}}{w_{i}}\geq\frac{T_{1}}{2},0\leq n_{i}\leq T_{1},\forall i\in[k],w\in\Delta_{\varepsilon_{1}/(8k)}(k)\bigg{\}},$
@@
one can see from (75) that, with probability exceeding $1-\delta/4$,
@@ B.2 Proof of Lemma 2
Before proceeding, let us introduce some additional notation. Let $\delta^{\prime}\coloneqq\frac{\delta}{4(T+k+1)}$, and recall
$\mathsf{OPT}=\min_{\pi\in\Delta(\mathcal{H})}\max_{1\leq i\leq k}L(h_{\pi},\mathrm{e}^{\mathsf{basis}}_{i})=\max_{w\in\Delta(k)}\min_{h\in\mathcal{H}}L(w,h)$
@@
$v^{t}\leq\min_{h\in\mathcal{H}}L(h,w^{t})-\mathsf{OPT}+\varepsilon_{1}=\min_{h\in\mathcal{H}}L(h,w^{t})-\max_{w\in\Delta(k)}\min_{h\in\mathcal{H}}L(h,w)+\varepsilon_{1}\leq\varepsilon_{1},\qquad\forall 1\leq t\leq T.$ (80)
@@
$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t+1}}{\sum_{i=1}^{k}W_{i}^{t}}\right)$ $\overset{\text{(i)}}{=}\log\left(\sum_{i=1}^{k}w_{i}^{t}\exp(\eta\widehat{r}^{t}_{i})\right)\overset{\text{(ii)}}{\leq}\log\left(\sum_{i=1}^{k}w_{i}^{t}\big{(}1+\eta\widehat{r}_{i}^{t}+\eta^{2}(\widehat{r}_{i}^{t})^{2}\big{)}\right)$